OpenStudy (anonymous):

The sum of three consecutive integers is 306. Find the integers. Write the three integers in order of smallest to largest:

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

a+(a+1)+(a+2)=306 3a+3=306 3a=303 a=101 101+102+103=306